# Dream to Manipulate: Compositional World Models Empowering Robot Imitation Learning with Imagination

### v1

@@ Abstract
A world model provides an agent with a representation of its environment, enabling it to predict the causal consequences of its actions. Current world models typically cannot directly and explicitly imitate the actual environment in front of a robot, often resulting in unrealistic behaviors and hallucinations that make them unsuitable for real-world applications. In this paper, we introduce a new paradigm for constructing world models that are explicit representations of the real world and its dynamics. By integrating cutting-edge advances in real-time photorealism with Gaussian Splatting and physics simulators, we propose the first compositional manipulation world model, which we call DreMa. DreMa replicates the observed world and its dynamics, allowing it to imagine novel configurations of objects and predict the future consequences of robot actions. We leverage this capability to generate new data for imitation learning by applying equivariant transformations to a small set of demonstrations. Our evaluations across various settings demonstrate significant improvements in both accuracy and robustness by incrementing actions and object distributions, reducing the data needed to learn a policy and improving the generalization of the agents. As a highlight, we show that a real Franka Emika Panda robot, powered by DreMa ’s imagination, can successfully learn novel physical tasks from just a single example per task variation (one-shot policy learning). Our project page and source code can be found in: https://leobarcellona.github.io/DreamToManipulate/
@@ 1 Introduction
Applying world models to real robots implies rendering realistic observations (Yang et al., 2023; Zhou et al., 2024) while simulating the dynamics (Li et al., 2024). Whereas most of the world models learn a statistical representation of the dynamics (Hafner et al., 2019; Gumbsch et al., 2024; Hansen et al., 2024; Yang et al., 2023), it is unrealistic to expect any training distribution to encompass all possible configurations of the world. This limits the current literature since agents need the ability to imagine beyond direct experience. We identify three requirements needed by an effective world model for a robotic agent. First, the world model must be inherently compositional, allowing the generation of novel and valid combinations of learned concepts (Zhou et al., 2024). Second, it must be object-centric, as the world is primarily composed of objects of varying number, type, and configuration (Spelke, 1990; Kipf et al., 2019). Third, manipulating the model should be simple to ensure that actions and consequences can be controllably modelled (Li et al., 2024).
@@
In this work, we propose a compositional manipulation world model, dubbed DreMa, that allows robotic agents to exploit physically reliable predictions to generalize to novel situations. Thus, it is an interactable physics-constrained representation that bridges robot imagination with real-world demonstrations. We shape DreMa by reconstructing the scene with object-centric Gaussian Splatting (Kerbl et al., 2023) and by embedding the obtained compositional representation into a physics simulator, e.g., PyBullet (Coumans and Bai, 2016). Gaussian Splatting reconstructs 3D scenes as a set of independent Gaussians that can be optimized and rendered, allowing for extremely high resolution and real-time graphics rendering. We infer compositional properties of rendered Gaussian Splats using foundational models (Ravi et al., 2024; Ren et al., 2024), allowing the agent to manipulate the object-centric model to create new scenes.
@@
Through DreMa, it is possible to execute robot policies directly in imagination while rendering realistic observations. This capability benefits imitation learning, whose objective is to solve tasks given example demonstrations (James et al., 2022). To ensure a reliable generalization of the learned policy, demonstrations should contain sufficient variations, which is impractical and time-consuming with real applications (Mandlekar et al., 2023). We show that by imposing equivariant transformations, namely roto-translation of the demonstrated actions and the objects, DreMa produces new high-quality data inside the world model and reduces the required demonstrations. We exploit imagination to generate new demonstrations that augment the original ones, while also using the proposed world model to verify that the new actions result in successful task execution and rendering of realistic observations.
@@
Our contributions are: (1) We are the first to propose a compositional manipulation world model, DreMa, that empowers robots with the ability to imagine in a grounded manner; (2) We propose to exploit the compositional structure of the world model to generate novel, valid, and realistic training demonstrations for imitation learning; (3) We empirically demonstrate in simulations and with a real Franka Emika Panda that DreMa improves agent generalizability even with a minimal number of demonstrations (+9.1% in simulation and +33.3% on real-world tasks).
@@ 2.1 World Models
World models were originally conceived to improve policy learning in reinforcement learning by predicting the consequences of actions on the environment (Ha and Schmidhuber, 2018; Hafner et al., 2019; 2020; 2023). State-of-the-art approaches often leverage generative models to build these world models (Jia et al., 2023; Hu et al., 2023; Yang et al., 2023; Bruce et al., 2024). While initially applied to game simulations (Ha and Schmidhuber, 2018; Hafner et al., 2019) and navigation tasks (Hu et al., 2023; Jia et al., 2023), world models are now being extended to robotics (Yang et al., 2023; Wu et al., 2023; Zhou et al., 2024). For example, UniSim (Yang et al., 2023) uses diffusion to generate possible future frames conditioned on robot actions. DayDreamer (Wu et al., 2023) demonstrates that skills like locomotion and manipulation can be learned using Dreamer (Hafner et al., 2019), while RoboDreamer (Zhou et al., 2024) uses compositionality to generate novel actions. The main limitation of these approaches is their inability to comprehensively capture the motion dynamics and physics specific to the robot’s environment, particularly when faced with scenarios outside their training domain (Yang et al., 2023). By contrast, we propose a compositional manipulation world model that integrates cutting-edge photorealism with Gaussian Splats and realistic dynamics with physics engines, showing we can tailor them for efficient imitation learning in real-world scenarios.
@@ 2.2 Simulation-based reconstructions
Thanks to recent advances in reconstruction (Kerbl et al., 2023) and in physics simulations (Coumans and Bai, 2016; Li et al., 2019), in the last year many propose to combine them both for realistic simulations of actual environments (Xie et al., 2024; Jiang et al., 2024; Meyer et al., 2024; Lou et al., 2024; Abou-Chakra et al., 2024; Ruan et al., 2024; Torne et al., 2024). Particle-based simulators are capable of modeling more complex dynamics than traditional rigid body simulators (Li et al., 2019), making them beneficial for robotics (Chen et al., 2024a). Recently, neural radiance fields have been applied to implicitly learn dynamic systems (Li et al., 2023; Whitney et al., 2024). The introduction of Gaussian Splatting (Kerbl et al., 2023) has opened avenues for novel approaches that bridge these simulation techniques with high-quality rendering methods (Xie et al., 2024; Jiang et al., 2024). However, they are not compositional (Xie et al., 2024) and object-centric (Xie et al., 2024; Jiang et al., 2024). PEGASUS (Meyer et al., 2024) combines Gaussian representations with mesh simulations to improve the 6D pose estimation, while Abou-Chakra et al. (2024) and Lou et al. (2024) show that Gaussian representations can be adapted to track moving objects. Both approaches do not allow agents to predict future states, rather focusing more on tracking in observed past trajectories. By contrast, we empirically demonstrate that our approach predicts useful estimation of future states. While simulating a replica of the environment can benefit policy evaluation (Li et al., 2024), two recent studies showed that scene reconstruction can enhance robot policies (Ruan et al., 2024; Torne et al., 2024). Ruan et al. (2024) reconstruct the environment to deform robot trajectories, while our object-centric approach focuses on the interaction between objects and the agent. While being close to our approach, Torne et al. (2024) requires to manually model the object-centric representation and needs adaptation between the simulation and the real world. Instead, we leverage the zero-shot capabilities of foundational models (Kirillov et al., 2023; Ren et al., 2024; Cheng et al., 2023) to automatically create a compositional world model and show the effectiveness of using transformed demonstrations for task learning without requiring extensive interaction with the world model.
@@ 2.3 Imitation learning and data augmentation
Figure 1: Overview of imagination with DreMa. From images of the environment, DreMa constructs a compositional manipulation world model using object-centric Gaussian Splatting. The world model is used to imagine novel demonstrations by transforming real demonstrations.
Imitation learning involves acquiring task knowledge from a dataset of demonstrations (James et al., 2022). Vision-based imitation learning focus on deriving policies directly from RGB or RGB-D images (Zeng et al., 2021; Shridhar et al., 2023; 2022; Goyal et al., 2024). PerAct (Shridhar et al., 2023), for instance, voxelizes the environment to infer actions in 3D. However, despite advancements in generalization, many approaches (Ze et al., 2023; Goyal et al., 2024) still require hundreds of demonstrations to learn robust policies. MIRA (Lin et al., 2023) employs NeRF’s novel view synthesis to simulate new camera views of the environment, while  Lu et al. (2025) applied Gaussian Splatting for action inference. In contrast, our approach leverages Gaussian Splatting to generate not only novel views but also new environmental configurations, enhancing generalization.
Data augmentation is an established technique (Krizhevsky et al., 2012; He et al., 2016). However, its application in robotics is challenging, as altering robot actions can unpredictably affect environments and tasks (Pitis et al., 2022). Despite this, Laskin et al. (2020) demonstrated that even simple augmentations can improve robot policies. Approaches such as generating counterfactual transitions (Pitis et al., 2022) and applying invariant transformations (Corrado and Hanna, 2024) enhance learning by expanding training datasets. However, these methods primarily recombine existing trajectories rather than generating entirely new configurations. In contrast, our approach imagines novel equivariant configurations of objects and actions, enabling more diverse and effective data augmentation. A related method, MimicGen (Mandlekar et al., 2023), uses object-centric segments for compositional data generation but requires multiple demonstrations. Our method achieves comparable benefits while working effectively with as few as one demonstration. Whereas some authors argue that imagining various object configurations is unrealistic (Zeng et al., 2021), we demonstrate that this is achievable in real-world scenarios.
@@ 3 Dream-to-Manipulate with Gaussian Splats & Physics Engines
In the following, we first discuss our compositional 3D scene representation based on Gaussian Splats, which forms the basis for our world model. After, we detail how we obtain each component of the world model given our 3D scene representation.
@@ 3.1 Representing 3D Objects with Gaussian Splatting
Central to the compositional manipulation world model is the inverse graphics function $\mathcal{G}=h(\mathbf{X})$ using Gaussian Splatting (Kerbl et al., 2023). The inverse graphics function returns a 3D representation of the world $\mathcal{G}$ given a sequence of $n$ images $\textbf{X}=(x_{1},...,x_{n})$, collected as a video (Kerbl et al., 2023). The key idea behind Gaussian Splatting is to ground many Gaussian blobs $g=\big{(}p,r,s,\alpha,c\big{)},\forall g\in\mathcal{G}$ physically in the 3D space. Per Gaussian, we have $p=[p_{x},p_{y},p_{z}]$ as its center , its orientation $r$ and scale $s=\big{(}s_{u},s_{v}\big{)}$ (the covariance), the encoded color $c$, and the opacity $\alpha$ of the color. That is, given any arbitrary 3D location and angle in space (corresponding to a 2D camera plane), Gaussian Splats can render a projected 2D image that resembles the true observed object (or scene). For rendering the appearance onto the 2D image, one simply $\alpha$-blends the colors of the Gaussians along each of the rays that connect the 3D point to the points on the 2D projected image. The Gaussian Splatting algorithm in the end optimizes the Gaussian parameters in $g$ to render the objects and the scene photorealistically for both known and unknown 3D locations (named novel view synthesis). Gaussian Splats have taken the graphics community by storm (Huang et al., 2024; Xie et al., 2024; Meyer et al., 2024) not only for their exceptional photorealism but also their real-time computation. For further details, we refer to the seminal work of Kerbl et al. (2023).
@@
Gaussian Splats generate photorealistic renderings of the 3D scene that are unstructured in terms of object semantics. Instead, we are interested in object-centric representations. To obtain object-centric Gaussian Splats, together with the input 2D images in $\mathbf{X}$ for our $K$ objects in our world, we introduce sets of segmentation masks, $\mathbf{Y}_{k}\in\mathcal{Y},k=1,\dots,K$. We consider $k$ to be varying so that not constrain in advance the number of objects in our world. Each object-set of segmentations $\mathbf{Y}_{k}=\big{(}y_{1,k},...,y_{n,k}\big{)}$ contains the 2D masks for the $k$-th object in all $N$ images in $\mathbf{X}$. With segmentation masks $\mathbf{Y}_{k}=\big{(}y_{1,k},...,y_{n,k}\big{)}$ for each object $k$ and for all $n$ frames, we can simply mask the respective video frames, $y_{n,k}\odot x_{n,k}$, to obtain the object-centric instances, that we can then feed to the Gaussian Splatting. Since in our setting the robot is also equipped with depth sensors, we further strengthen the original Gaussian Splatting optimization by including the depth measurements $x^{\text{depth}}\in\mathbf{X}_{\text{depth}}$ for additional supervision, $\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda_{\text{n}}\mathcal{L}_{\text{n}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}$.Here $\mathcal{L}_{\text{rec}}$ and $\mathcal{L}_{n}$ are the reconstruction and normal losses defined in the original 2DGS. The extra depth regularization, $\mathcal{L}_{\text{depth}}=\|\hat{x}_{GS}^{\text{depth}}-x^{\text{depth}}\|_{1}$ , uses the $\ell_{1}$ norm to penalize Gaussians that have outlier depth. We set the hyperparameters $\lambda_{\text{normal}}$ and $\lambda_{\text{depth}}$ for all experiments empirically as in Huang et al. (2024). Our object-centric inverse graphics function takes the form $\mathcal{G}_{k}=h\big{(}\mathbf{X},\mathbf{X}_{\text{depth}},\mathbf{Y}_{k}\big{)}$.
@@
It is important to note that since the standard Gaussian Splats only work with (static) scenes, we cannot account for deformable or articulated objects. An exciting future direction is to bootstrap on the significant progress in Dynamic Gaussian Splatting (Luiten et al., 2023) to model also articulation and deformation in the objects of our world.
@@
3.2 Decomposing the Scene to Object-centric Regions
Decomposing the high-dimensional scene representation into objects (Locatello et al., 2020; Seitzer et al., 2023; Wu et al., 2024) that are independently controllable greatly simplifies the control, and the manipulation of the dynamics in the environment (Zadaianchuk et al., 2021; 2022). Inspired by recent work on embeddable 3D assets (Deitke et al., 2023; Wu et al., 2024), we ideally want to learn Gaussian Splats that belong to the same object. We take advantage of the zero-shot capabilities of open-vocabulary tracking models (Cheng et al., 2023; Wang et al., 2023; Ravi et al., 2024) to ensure that we obtain consistent predictions. Our compositional world model uses the DEVA open-vocabulary tracker to extract the objects and their segmentation masks $\mathcal{Y}_{k}$ in the video $\mathcal{X}$ of the scene. Since DEVA discovers objects given prompts, we use general and minimal prompts: object for all the objects and table for segmenting the table of our table-top robot. Having computed the segmentation masks $\mathcal{Y}_{k}=\big{(}y_{1,k},...,y_{n,k}\big{)}$ object $k$ and for all $n$ frames, we can learn object-centric Gaussian Splats as described in Sec. 3.1. Quite naturally, advances in the object-centric (Didolkar et al., 2024; Zadaianchuk et al., 2024; Wu et al., 2024) and open-vocabulary (Ravi et al., 2024; Bianchi et al., 2024) literature can be seamlessly integrated into our compositional manipulation world model.

### v2

@@ Abstract
A world model provides an agent with a representation of its environment, enabling it to predict the causal consequences of its actions. Current world models typically cannot directly and explicitly imitate the actual environment in front of a robot, often resulting in unrealistic behaviors and hallucinations that make them unsuitable for real-world robotics applications. To overcome those challenges, we propose to rethink robot world models as learnable digital twins. We introduce DreMa, a new approach for constructing digital twins automatically using learned explicit representations of the real world and its dynamics, bridging the gap between traditional digital twins and world models. DreMa replicates the observed world and its structure by integrating Gaussian Splatting and physics simulators, allowing robots to imagine novel configurations of objects and to predict the future consequences of robot actions thanks to its compositionality. We leverage this capability to generate new data for imitation learning by applying equivariant transformations to a small set of demonstrations. Our evaluations across various settings demonstrate significant improvements in accuracy and robustness by incrementing actions and object distributions, reducing the data needed to learn a policy and improving the generalization of the agents. As a highlight, we show that a real Franka Emika Panda robot, powered by DreMa’s imagination, can successfully learn novel physical tasks from just a single example per task variation (one-shot policy learning). Our project page can be found in: https://dreamtomanipulate.github.io/.
@@ 1 Introduction
Applying world models to real robots implies rendering realistic observations (Yang et al., 2023; Zhou et al., 2024) while simulating the dynamics (Li et al., 2024). Whereas most of the world models learn a statistical representation of the dynamics (Hafner et al., 2019; Gumbsch et al., 2024; Hansen et al., 2024; Yang et al., 2023), it is unrealistic to expect any training distribution to encompass all possible configurations of the world. This limits the current literature, as agents need the ability to imagine beyond direct experience. We identify three requirements needed by an effective world model for a robotic agent. First, it must be inherently compositional, allowing the generation of novel and valid combinations of learned concepts (Zhou et al., 2024). Second, it must be object-centric, as the world is primarily composed of objects of varying number, type, and configuration (Spelke, 1990; Kipf et al., 2019). Third, manipulating the model should be simple to ensure that actions and consequences can be controllably modelled (Li et al., 2024).
@@
In this work, we propose a compositional manipulation world model, named DreMa, that allows robotic agents to exploit physically reliable predictions to generalize to novel situations. Thus, it is an interactable physics-constrained representation that bridges robot imagination with real-world demonstrations following recent advances in learnable digital twins and explicit world models, which both enable agents to simulate and reason about their environment (discussed in App. E). We shape DreMa by reconstructing the scene with object-centric Gaussian Splatting (Kerbl et al., 2023) and by embedding the obtained compositional representation into a physics simulator, e.g., PyBullet (Coumans and Bai, 2016). Gaussian Splatting reconstructs 3D scenes as a set of independent Gaussians that can be optimized and rendered, allowing for extremely high resolution and real-time graphics rendering. We infer compositional properties of rendered Gaussian Splats using foundational models (Ravi et al., 2024; Ren et al., 2024), allowing the agent to manipulate the object-centric model to create new scenes.
@@
Through DreMa, it is possible to execute robot policies directly in imagination while rendering realistic observations. This capability benefits imitation learning, whose objective is to solve tasks given example demonstrations (James et al., 2022). To ensure a reliable generalization of the learned policy, demonstrations should contain sufficient variations, which is impractical and time-consuming with real applications (Mandlekar et al., 2023). We show that imposing equivariant transformations, namely roto-translation of the demonstrated actions and the objects, DreMa produces new high-quality data inside the world model and reduces the required demonstrations. We exploit imagination to generate new demonstrations that augment the original ones, while also using the proposed world model to verify that the new actions result in successful task execution and rendering of realistic observations.
@@
Our contributions are: (1) We are the first to propose a compositional manipulation world model, DreMa, that empowers robots with the ability to imagine in a grounded manner; (2) We propose to exploit the compositional structure of the world model to generate novel, valid, and realistic training demonstrations for imitation learning; (3) We empirically demonstrate in simulations and with real robots that DreMa improves agent generalizability even with a minimal number of demonstrations.
Figure 1: Overview of imagination with DreMa, which builds a compositional manipulation world model from environment images using object-centric Gaussian Splatting to generate novel demonstrations by transforming real ones.
@@ 2.1 World Models
World models were originally conceived to improve policy learning in reinforcement learning by predicting the consequences of actions on the environment (Ha and Schmidhuber, 2018; Hafner et al., 2019; 2020; 2023). State-of-the-art approaches often leverage generative models to build these world models (Jia et al., 2023; Hu et al., 2023; Yang et al., 2023; Bruce et al., 2024). While initially applied to game simulations (Ha and Schmidhuber, 2018; Hafner et al., 2019) and navigation tasks (Hu et al., 2023; Jia et al., 2023), world models are now being extended to robotics (Yang et al., 2023; Wu et al., 2023; Zhou et al., 2024). For example, UniSim (Yang et al., 2023) uses diffusion to generate possible future frames conditioned on robot actions. DayDreamer (Wu et al., 2023) demonstrates that skills such as locomotion and manipulation can be learned using Dreamer (Hafner et al., 2019), while RoboDreamer (Zhou et al., 2024) uses compositionality to generate novel actions. The main limitation of these approaches is their inability to comprehensively capture the motion dynamics and physics specific to the robot’s environment, particularly when faced with scenarios outside their training domain (Yang et al., 2023). By contrast, we propose a compositional manipulation world model that integrates cutting-edge photorealism with Gaussian Splats and realistic dynamics with physics engines, showing we can tailor them for efficient imitation learning in real-world scenarios.
@@ 2.2 Simulation-based reconstructions
Thanks to recent advances in reconstruction (Kerbl et al., 2023) and in physics simulations (Coumans and Bai, 2016; Li et al., 2019), in the last year many have proposed to combine both for realistic simulations of actual environments (Xie et al., 2024; Jiang et al., 2024; Meyer et al., 2024; Lou et al., 2024; Abou-Chakra et al., 2024; Ruan et al., 2024; Torne et al., 2024). Particle-based simulators are capable of modeling more complex dynamics than traditional rigid body simulators (Li et al., 2019), making them beneficial for robotics (Chen et al., 2024a). Recently, neural radiance fields have been applied to implicitly learn dynamic systems (Li et al., 2023; Whitney et al., 2024). The introduction of Gaussian Splatting (Kerbl et al., 2023) has opened avenues for novel approaches that bridge these simulation techniques with high-quality rendering methods (Xie et al., 2024; Jiang et al., 2024). However, they are not compositional (Xie et al., 2024) or object-centric (Xie et al., 2024; Jiang et al., 2024). PEGASUS (Meyer et al., 2024) combines Gaussian representations with mesh simulations to improve the 6D pose estimation, while Abou-Chakra et al. (2024) and Lou et al. (2024) show that Gaussian representations can be adapted to track moving objects. Both approaches do not allow agents to predict future states, rather focusing more on tracking in observed past trajectories. In contrast, we empirically demonstrate that our approach predicts useful estimation of future states. While simulating a replica of the environment can benefit policy evaluation (Li et al., 2024), two recent studies showed that scene reconstruction can enhance robot policies (Ruan et al., 2024; Torne et al., 2024). Ruan et al. (2024) reconstruct the environment to deform robot trajectories, while our object-centric approach focuses on the interaction between objects and the agent. Although being close to our approach, Torne et al. (2024) requires to manually model the object-centric representation and needs adaptation between the simulation and the real world. Instead, we leverage the zero-shot capabilities of foundational models (Kirillov et al., 2023; Ren et al., 2024; Cheng et al., 2023) to automatically create a compositional world model and show the effectiveness of using transformed demonstrations for task learning without requiring extensive interaction with the world model.
@@ 2.3 Imitation learning and data augmentation
Imitation learning involves acquiring task knowledge from a dataset of demonstrations (James et al., 2022). Vision-based imitation learning focuses on deriving policies directly from RGB or RGB-D images (Zeng et al., 2021; Shridhar et al., 2023; 2022; Goyal et al., 2024). PerAct (Shridhar et al., 2023), for instance, voxelizes the environment to infer actions in 3D. However, despite advancements in generalization, many approaches (Ze et al., 2023; Goyal et al., 2024) still require hundreds of demonstrations to learn robust policies. MIRA (Lin et al., 2023) employs NeRF’s novel view synthesis to simulate new camera views of the environment, while  Lu et al. (2025) applied Gaussian Splatting for action inference. In contrast, our approach leverages Gaussian Splatting to generate not only novel views but also new environmental configurations, enhancing generalization.
Data augmentation is an established technique (Krizhevsky et al., 2012; He et al., 2016). However, its application in robotics is challenging, as altering robot actions can unpredictably affect environments and tasks (Pitis et al., 2022). Despite this, Laskin et al. (2020) demonstrated that even simple augmentations can improve robot policies. Approaches such as generating counterfactual transitions (Pitis et al., 2022) and applying invariant transformations (Corrado and Hanna, 2024) enhance learning by expanding training datasets. However, these methods primarily recombine existing trajectories rather than generate entirely new configurations. In contrast, our approach imagines novel equivariant configurations of objects and actions, enabling more diverse and effective data augmentation. A related method, MimicGen (Mandlekar et al., 2023), uses object-centric segments for compositional data generation but requires multiple demonstrations. Our method achieves comparable benefits while working effectively with as few as one demonstration. Whereas some authors argue that imagining various object configurations is unrealistic (Zeng et al., 2021), we demonstrate that this is achievable in real-world scenarios.
@@ 3 Dream-to-Manipulate with Gaussian Splats & Physics Engines
In the following, we first discuss our compositional 3D scene representation based on Gaussian Splats, which forms the basis for our world model. Afterward, we detail how we obtain each component of the world model given our 3D scene representation.
@@ 3.1 Representing 3D Objects with Gaussian Splatting
Central to the compositional manipulation world model is the inverse graphics function $\mathcal{G}=h(\mathbf{X})$ using Gaussian Splatting (Kerbl et al., 2023). The inverse graphics function returns a 3D representation of the world $\mathcal{G}$ given a sequence of $n$ images $\textbf{X}=(x_{1},...,x_{n})$, collected as a video (Kerbl et al., 2023). The key idea behind Gaussian Splatting is to ground many Gaussian blobs $g=\big{(}p,r,s,\alpha,c\big{)},\forall g\in\mathcal{G}$ physically in the 3D space. Per Gaussian, we have $p=[p_{x},p_{y},p_{z}]$ as its center , its orientation $r$ and scale $s=\big{(}s_{u},s_{v}\big{)}$ (the covariance), the encoded color $c$, and the opacity $\alpha$ of the color. That is, given any arbitrary 3D location and angle in space (corresponding to a 2D camera plane), Gaussian Splats can render a projected 2D image that resembles the true observed object (or scene). For rendering the appearance onto the 2D image, one simply $\alpha$-blends the colors of the Gaussians along each of the rays that connect the 3D point to the points on the 2D projected image. In the end, the Gaussian Splatting algorithm optimizes the Gaussian parameters in $g$ to render the objects and the scene photorealistically for both known and unknown 3D locations (named novel view synthesis). Gaussian Splats have taken the graphics community by storm (Huang et al., 2024; Xie et al., 2024; Meyer et al., 2024) not only for their exceptional photorealism but also for their real-time computation. For further details, we refer to the seminal work of Kerbl et al. (2023).
@@
Gaussian Splats generate photorealistic renderings of the 3D scene that are unstructured in terms of object semantics. Instead, we are interested in object-centric representations. To obtain object-centric Gaussian Splats, together with the input 2D images in $\mathbf{X}$ for our $K$ objects in our world, we introduce sets of segmentation masks, $\mathbf{Y}_{k}\in\mathcal{Y},k=1,\dots,K$. We consider $k$ to be varying so that not constrain in advance the number of objects in our world. Each object-set of segmentations $\mathbf{Y}_{k}=\big{(}y_{1,k},...,y_{n,k}\big{)}$ contains the 2D masks for the $k$-th object in all $N$ images in $\mathbf{X}$. With segmentation masks $\mathbf{Y}_{k}=\big{(}y_{1,k},...,y_{n,k}\big{)}$ for each object $k$ and for all $n$ frames, we can simply mask the respective video frames, $y_{n,k}\odot x_{n,k}$, to obtain the object-centric instances, which we can then feed to the Gaussian Splatting. Since in our setting the robot is also equipped with depth sensors, we further strengthen the original Gaussian Splatting optimization by including the depth measurements $x^{\text{depth}}\in\mathbf{X}_{\text{depth}}$ for additional supervision, $\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda_{\text{n}}\mathcal{L}_{\text{n}}+\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}$.Here $\mathcal{L}_{\text{rec}}$ and $\mathcal{L}_{n}$ are the reconstruction and normal losses defined in the original 2DGS. The extra depth regularization, $\mathcal{L}_{\text{depth}}=\|\hat{x}_{GS}^{\text{depth}}-x^{\text{depth}}\|_{1}$ , uses the $\ell_{1}$ norm to penalize Gaussians that have outlier depth. We set the hyperparameters $\lambda_{\text{normal}}$ and $\lambda_{\text{depth}}$ for all experiments empirically as in Huang et al. (2024). Our object-centric inverse graphics function takes the form $\mathcal{G}_{k}=h\big{(}\mathbf{X},\mathbf{X}_{\text{depth}},\mathbf{Y}_{k}\big{)}$.
@@
It is important to note that since the standard Gaussian Splats only work with (static) scenes, we cannot account for deformable or articulated objects. An exciting future direction is to bootstrap on the significant progress in Dynamic Gaussian Splatting (Luiten et al., 2023) to automatically model also articulation (see App. G) and deformation.
@@
3.2 Decomposing the Scene into Object-centric Regions
Decomposing the high-dimensional scene representation into objects (Locatello et al., 2020; Seitzer et al., 2023; Wu et al., 2024) that are independently controllable greatly simplifies control and the manipulation of the dynamics in the environment (Zadaianchuk et al., 2021; 2022). Inspired by recent work on embeddable 3D assets (Deitke et al., 2023; Wu et al., 2024), we ideally want to learn Gaussian Splats that belong to the same object. We take advantage of the zero-shot capabilities of open-vocabulary tracking models (Cheng et al., 2023; Wang et al., 2023; Ravi et al., 2024) to ensure that we obtain consistent predictions. Our compositional world model uses the DEVA open-vocabulary tracker to extract the objects and their segmentation masks $\mathcal{Y}_{k}$ in the video $\mathcal{X}$ of the scene. Since DEVA discovers objects given prompts, we use general and minimal prompts: object for all the objects and table for segmenting the table of our table-top robot. Having computed the segmentation masks $\mathcal{Y}_{k}=\big{(}y_{1,k},...,y_{n,k}\big{)}$ object $k$ and for all $n$ frames, we can learn object-centric Gaussian Splats as described in Sec. 3.1. Advances in the object-centric (Didolkar et al., 2024; Zadaianchuk et al., 2024; Wu et al., 2024) and open-vocabulary (Ravi et al., 2024; Bianchi et al., 2024) literature can be seamlessly integrated into our compositional manipulation world model.